# Leveraging Large Language Models for Suicide Detection on Social Media with Limited Labels

### v1

@@ III-A Deep learning approaches for text classification
Feature extraction approaches begin with extracting features from the input text and then training a classifier for the downstream task. Some traditional methods such as Word2Vec [6] and GloVe [7] have been widely used. These methods represent words as high-dimensional vectors learned from word co-occurrence statistics. However, they often struggle to capture the complexities of language as they cannot capture the context in the input. Transformer-based models (e.g., BERT [8], LLaMA3 [4], GPT-4 [9]), have recently become more commonly used as feature extractors due to their high performance [10]. These models, which were trained on large corpora of text, use attention mechanisms [11] to capture the context of words in a sentence. This enables them to understand the meaning of words in a specific context. The embeddings produced by the LLM’s pretrained parameters serve as input features for training traditional machine learning classifiers like XGBoost [12]. By freezing the LLM’s parameters during this process, we can leverage its ability to capture intricate linguistic patterns without the need for extensive fine-tuning. This hybrid approach combines the strengths of LLMs in feature extraction with the simplicity of traditional classifiers, offering an effective and efficient method for text classification.
@@ IV-A Generating pseudo-labels for unlabeled data
We employ Qwen2-72B-Instruct  [^1] https://huggingface.co/Qwen/Qwen2-72B-Instruct to generate pseudo-label for the unlabeled posts, as it is one of the performant open-source models with an extended context window of up to $128K$ tokens [3]. We use greedy-decoding (i.e., choosing the highest probability output at each step of the sequence generation) and set the maximum length of the generated tokens to $1024$. We also experiment with other LLMs, such as LLaMA3 [4] and Mistral [37, 38] family models, refer to Table VII and Section VI for further discussion.
@@ V Results
Figure 4: Comparison of F1 Scores for our models on the Public Board Test Set. The ensemble model shows its robustness and significantly outperforms the individual models, demonstrating an improvement of approximately 5% points in the F1 Score on the new test set.
@@
Fig. 4 presents the comparison of F1 Scores for our models on the test set from the Public Board of the competition. It is noticeable that the F1 scores in the test set are lower than those in the validation set. This discrepancy may be attributed to the differences in label distribution between the two datasets. That said, the ensemble model demonstrates strong robustness and still performs well on the test set. Specifically, the ensemble model achieves the highest F1 score on the test set at $0.770$, significantly better than any individual model by $5$% points.
@@ VI Analysis and Discussion
Among $47$ registered teams, our solution achieved a score of $75.436$ the highest score on overall evaluation, based on model performance ($6$-th on the private board with an F1 score of $0.7312$), approach innovation, and report quality.
@@ VII Conclusion
However, our work has some limitations. The ensemble model, particularly Qwen2-72B-Instruct, requires a substantial amount of time for inference due to its large size, which poses challenges for real-time deployment. To mitigate this, smaller distillation models [44] could offer a more practical solution. Additionally, enhancing suicide detection using LLMs would benefit from better prompt engineering by domain experts. Breaking prompts into smaller, more detailed questions can help guide LLMs more effectively and interoperably. Another promising direction for future research is the integration of visual data (i.e., images and video), commonly found in user posts, which may require multimodal LLMs.

### v2

@@ III-A Deep learning approaches for text classification
Feature extraction approaches begin with extracting features from the input text and then training a classifier for the downstream task. Some traditional methods such as Word2Vec [6] and GloVe [7] have been widely used. These methods represent words as high-dimensional vectors learned from word co-occurrence statistics. However, they often struggle to capture the complexities of language as they cannot capture the context in the input. Transformer-based models (e.g., BERT [8], LLaMA3 [4], GPT-4 [9]), have recently become more commonly used as feature extractors due to their high performance [10]. These models, which were trained on large corpora of text, use attention mechanisms [11] to capture the context of words in a sentence. This enables them to understand the meaning of words in a specific context. The embeddings produced by the LLM’s pretrained parameters serve as input features for training traditional machine learning classifiers like XGBoost [12]. By freezing the LLM’s parameters during this process, we can leverage its ability to capture intricate linguistic patterns without the need for extensive fine-tuning. This hybrid approach combines the strengths of LLMs in feature extraction with the simplicity of traditional classifiers, offering an efficient method for text classification.
@@ IV-A Generating pseudo-labels for unlabeled data
We employ Qwen2-72B-Instruct  [^1] https://huggingface.co/Qwen/Qwen2-72B-Instruct to generate pseudo-label for the unlabeled posts, as it is one of the performant open-source models with an extended context window of up to $128K$ tokens [3]. We use greedy-decoding (i.e., choosing the highest probability output at each step of the sequence generation) and set the maximum length of the generated tokens to $1024$. We also experiment with other LLMs, such as LLaMA3 [4] and Mistral family models [37, 38], refer to Table VII and Section VI for further discussion.
@@ V Results
Figure 4: Comparison of F1 Scores for our models on the Public Board Test Set containing $\mathbf{100}$ posts. The ensemble model shows its robustness and significantly outperforms the individual models, demonstrating an improvement of approximately $5$% points in the F1 Score on the new test set.
@@
Fig. 4 presents the comparison of F1 Scores for our models on the test set from the Public Board of the competition. It is noticeable that the F1 scores in the test set which consists of $100$ new user posts, are lower than those in the validation set. This discrepancy may be attributed to the differences in label distribution between the two datasets. That said, the ensemble model demonstrates strong robustness and still performs well on the test set. Specifically, the ensemble model achieves the highest F1 score on the test set at $0.770$, significantly better than any individual model by $5$% points.
@@ VI Analysis and Discussion
Among $47$ registered teams, our solution achieved a score of $75.436$ the highest score on overall evaluation, based on model performance ($6$-th on the private board containing a new set of $100$ user posts, with an F1 score of $0.7312$), approach innovation, and report quality.
@@ VII Conclusion
However, our work has some limitations. The ensemble model, particularly Qwen2-72B-Instruct, requires a substantial amount of time for inference due to its large size, which poses challenges for real-time deployment. To mitigate this, smaller distillation models [44] could offer a more practical solution. Additionally, enhancing suicide detection using LLMs would benefit from better prompt engineering by domain experts. Breaking prompts into smaller, more detailed questions can help guide LLMs more effectively and interoperably. Another promising direction for future research is integrating visual data, such as images and videos commonly found in user posts, which may necessitate the use of multimodal LLMs.